An educator believes that new reading activities for elementary school children will improve reading comprehension scores. She randomly assigns third graders to an eight-week program in which some will use these activities and others will experience traditional teaching methods. At the end of the experiment, both groups take a reading comprehension exam. Their Scores are shown below.

New Activities: 23, 34, 34, 34, 35, 37, 42, 43, 44, 46, 47, 48, 49, 55

Control: 12, 13, 22, 26, 29, 30, 32, 32, 32, 32, 32 34, 36, 43, 45, 45, 64

**This data was given in a stem and leaf display

Test an Appropriate hypothesis at the state your conclusion. Use a 5% significance level or \alpha =0.05.

t-statistic, t=2.13

The P-value is 0.021

1) Use the Wilcoxon rank sum test to test an appropriate hypothesis. Compute​ W, the test​ statistic, for the New Activities group.

W=?
2) Find the P-value.

The P-value is ?

**I actually have the answer to this but cannot wrap my head around how they got the answer. I ran into that troube to regarding the dataset not being paired but i literally am not given anything else.

Answer: W=288 and the P-value=0.006

HELP!
Answer #1

1.) As per Question, the Hypotheses becomes:

Ho: There is no change in average score of students who attended the program and who doesn't.

Mx = My

Ha: There is change in average score of students who attended the program and who doesn't.

Mx \neq My

Given data for New Activities:

New(X)
23
34
34
34
35
37
42
43
44
46
47
48
49
55

Mean, Mx = 40.7857

New(X) Diff (X - Mx) Sq. Diff (X - Mx)2
23 -17.786 316.332
34 -6.786 46.046
34 -6.786 46.046
34 -6.786 46.046
35 -5.786 33.474
37 -3.786 14.332
42 1.214 1.474
43 2.214 4.903
44 3.214 10.332
46 5.214 27.189
47 6.214 38.617
48 7.214 52.046
49 8.214 67.474
55 14.214 202.046

Nx: 14
dfx = N - 1 = 14 - 1 = 13
Mx: 40.79
Sum Of Square(SSx): 906.357
s2x = SSx/(Nx - 1) = 906.357/(14-1) = 69.72

Given Control Data:

Control(Y)
12
13
22
26
29
30
32
32
32
32
32
34
36
43
45
45
64

My = 32.8823

Control(Y) Diff (Y - My) Sq. Diff (Y - My)2
12 -20.882 436.073
13 -19.882 395.308
22 -10.882 118.426
26 -6.882 47.367
29 -3.882 15.073
30 -2.882 8.308
32 -0.882 0.779
32 -0.882 0.779
32 -0.882 0.779
32 -0.882 0.779
32 -0.882 0.779
34 1.118 1.249
36 3.118 9.720
43 10.118 102.367
45 12.118 146.837
45 12.118 146.837
64 31.118 968.308

Ny: 17
dfy = N - 1 = 17 - 1 = 16
My: 32.8823
SSy: 2399.765
s2y = SSy/(Ny - 1) = 2399.765/(17-1) = 149.99

Formula for t-statistic is:

t_independent_means.png

T-value Calculation

s2p = ((dfx/(dfx + dfy)) * s2x) + ((dfy/(dfx + dfy)) * s2y)

s2p = ((13/29) * 69.72) + ((16/29) * 149.99) = 114

s2Mx = s2p/Nx = 114/14 = 8.1433
s2My = s2p/Ny = 114/17 = 6.7062

t = (Mx- My)/√(s2Mx + s2My) = 7.9033/√14.85 = 2.05

So, t = 2.05

Hence corresponding p -value is, p = 0.0247

Since, given alpha = 0.05

P < 0.05, so we can reject the Null Hypotheses(Ho)

The alternate Hypotheses(Ha) is accepted.

2.)

For Wilcoxon rank sum test, Hypotheses are:

Ho: Median Scores for both the groups is same

MDx = MDy

Ha: Median Scores for both the groups is not same

MDx\neq MDy

First, we put both samples together and organize it in ascending order, which is shown in the table below:

Values Sample
12 2
13 2
22 2
23 1
26 2
29 2
30 2
32 2
32 2
32 2
32 2
32 2
34 2
34 1
34 1
34 1
35 1
36 2
37 1
42 1
43 2
43 1
44 1
45 2
45 2
46 1
47 1
48 1
49 1
55 1
64 2

Now, that the values that are in ascending order are assigned ranks to them, taking care of assigning the average rank to values with rank ties

Values Sample Rank Rank (Adjusted for ties)
12 2 1 1
13 2 2 2
22 2 3 3
23 1 4 4
26 2 5 5
29 2 6 6
30 2 7 7
32 2 8 10
32 2 9 10
32 2 10 10
32 2 11 10
32 2 12 10
34 2 13 14.5
34 1 14 14.5
34 1 15 14.5
34 1 16 14.5
35 1 17 17
36 2 18 18
37 1 19 19
42 1 20 20
43 2 21 21.5
43 1 22 21.5
44 1 23 23
45 2 24 24.5
45 2 25 25
46 1 26 26
47 1 27 27
48 1 28 28
49 1 29 29
55 1 30 30
64 2 31 31

Note : For Rank Adjusted with ties, we will take average of corresponding ranks and fill the same

average rank for all values

The sum of adjusted ranks for sample 1 is:

R1= 4+14.5+14.5+14.5+17+19+20+21.5+23+26+27+28+29+30 = 288

and the sum of ranks of sample 2 is:

R2 = 1+2+3+5+6+7+10+10+10+10+10+14.5+18+21.5+24.5+24.5+31 = 208

Hence, the test statistic is W = R1 = 288

To get P value, we need to calulate Z statistic:

The z-statistic is computed as follows:

z = [W - (n1* ​​(n1​+n2​+1) / 2) ] / \sqrt{[n1 * n2 * (n1 + n2 +1)] / 12}

z = 2.54

Checking for p value in z score tabe, we get

p = 0.0055 = 0.006

since p < 0.05, we can reject the Null Hypotheses.

So, Alternate Hypotheses is accepted and there is enough evidence for below expression

MDx\neq MDy